A researcher wishes to test the claim of a particular cereal manufacturer that the mean weight of cereal in the boxes is less than 300g. A sample of 50 boxes yields a sample mean weight of 296g. Assume that the population standard deviation is 5g.


Can we conclude that the claim is true? Test at α = 0.05.

Obtain a 95% confidence interval for μ.


1
Expert's answer
2020-08-25T11:00:19-0400

ANSWER 1)

Let us consider the null hypothesis H0 : "\mu = 300" against the claim of the researcher H1 : "\mu < 300" .


Test Criteria : Reject the null hypothesis H0 if Zcalc < Z"\alpha"


Given that,

Sample quantity N = 50

Sample mean U = 296

Standard deviation "\sigma = 5"

"\mu = 300"

We can find Zcalc = "\frac{U-\mu}{\frac{\sigma}{\sqrt N}}"

Zcalc = "\frac{296-300}{\frac{5}{\sqrt 50}} = -5.6568"


Since it is a left tailed test we will have to find the value of Z(1-"\alpha")


Z(1-"\alpha") = Z(1-0.05) = -1.6449


Hence, Zcalc < Z(1-"\alpha") as -5.6568 < -1.6449


So we can reject the null hypothesis and say that the claim of the researcher is true.


ANSWER 2)


The 95% confidence interval for a left-tailed test (when level of significance = "\alpha" ) is :

"(-\infin, U - Z_{1-\alpha}*\frac{\sigma}{\sqrt n}]"

"(-\infin, 296 - (-1.6449)*\frac{5}{\sqrt 50}]"

"(-\infin, 297.1631]"


95% confidence interval for μ is "(-\infin, 297.1631]"


But in this case since the population mean μ cannot be "-\infin" since it is in terms of weight so we can reduce the interval to [0, 297.1631]